OpenStudy (anonymous):

The position of a particle at time t is s(t)=(t^3) +t . How do I ESTIMATE the instantaneous velocity at t=1 without using a calculator?

OpenStudy (anonymous):

Well what do you know about the relationship[ between position and velocity?

OpenStudy (anonymous):

velocity is the derivative of the position?

OpenStudy (anonymous):

yes so you find s'(t) = v(t) thgen evaluate v(t) at t=1 so v(1) = ?

OpenStudy (anonymous):

oh, i'm not supposed to find the derivative because my teacher wants us to use: as the limit approaches...

OpenStudy (anonymous):

i mean, both methods are finding the derivative, but i'm not supposed to "know the easy way (your way)" yet

OpenStudy (anonymous):

ah so she wants you to use the definition of the derivative meaning.... \[f'(x) = \lim_{h \rightarrow 0} \frac{ f(x+h) - f(x) }{ h }\]

OpenStudy (anonymous):

yes, like that, except for the minor change of changing the 'f' to 's' and 'h' to 't'

OpenStudy (anonymous):

well x becomes t but f does become s yes so

OpenStudy (anonymous):

if i want the instantanoues velocity at t=1, should i write as t ->0 or as t->1 under 'lim'?

OpenStudy (anonymous):

nah the h stays the x is what becomes the t

OpenStudy (anonymous):

so this is what you are after \[v(t) = \lim_{h \rightarrow 0}\frac{ s(t+h) - s(t) }{ h }\]

OpenStudy (anonymous):

what does the 'h' refer to?

OpenStudy (anonymous):

it is basically the change in the variable so some textbooks refer to the h as this.... \[f'(x) = \lim_{\Delta x \rightarrow 0} \frac{ f(x+\Delta x )-f(x)}{ \Delta x }\]

OpenStudy (anonymous):

so it's always delta (variable) ->0 and not approaching another value, like 1?

OpenStudy (aum):

\[ v(1) = \lim_{\Delta t \rightarrow 0}\frac{ s(1+\Delta t) - s(1) }{ \Delta t } \]

OpenStudy (anonymous):

no it approaches 0 because if you were to use the formula the 0 makes all the change in disappear and you are left with a formula that you can plug your value in or you can shortcut it the way aum has by putting the 1 in right way

OpenStudy (anonymous):

um what do i do if i want an estimate of v(1)?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

your estimate is when you plug the value in. If it is asking for a estimate without putting a value in you can't really to be honest

OpenStudy (aum):

\[ v(1) = \lim_{h \rightarrow 0}\frac{ s(1+h) - s(1) }{ h } = \lim_{h \rightarrow 0}\frac{ (1+h)^3 + (1+h) - (1^3 + 1) }{ h } = ? \]

OpenStudy (anonymous):

yea that will be your answer

OpenStudy (aum):

\( (1+h)^3 = 1 + 3h + 3h^2 + h^3\)

OpenStudy (anonymous):

in the end i got 4+4h. is that correct?

OpenStudy (anonymous):

make the h become 0 you get 4

OpenStudy (aum):

\[ v(1) = \lim_{h \rightarrow 0} 4 + 4h = 4\]

OpenStudy (anonymous):

oh yeah! i forgot to plug in the 0! and the 4 here is the same as 4 i would have gotten if i used the shorter method, i.e. 3t squared + 1

OpenStudy (aum):

Yes. v(t) = 3t^2 + 1 v(1) = 3 + 1 = 4

OpenStudy (aum):

So yours is not really an estimate but an accurate value done through limits instead of using derivatives formula.

OpenStudy (anonymous):

As I said i wouldn't know how to do it as an estimate

OpenStudy (anonymous):

my textbook says "To estimate the instantaneous rate of change at x = x0, compute the average rate of change over several intervals[x0,x1](or[x1,x0]) for x1 close to x0. "

OpenStudy (anonymous):

ah they want you to do that?

OpenStudy (aum):

In that case, v(1) = ( s(2) - s(0) ) / ( 2 - 0 )

OpenStudy (anonymous):

and to find the estimated instantaenous velocity, "compute the average rate for several intervals lying to the left and right of" t=1

OpenStudy (anonymous):

hilbert, i think so - but i would need a calculator for the decimals (like t=1.0001 and t= .9999), but the teacher says i can't use a calculator

OpenStudy (anonymous):

I find this question flawed to be honest

OpenStudy (aum):

\[ v(1) \approx \frac { s(2) - s(0) }{ 2 - 0 } = \frac{2^3 + 2 - 0^3 - 0}{2-0} = \frac{10}{2} = 5 \]

OpenStudy (anonymous):

aum, i'm curious about why it's s(2)?

OpenStudy (anonymous):

he is taking a number to the right of 1

OpenStudy (aum):

one unit to the left of t = 1 and one unit to the right of t = 1 t = 0 and t = 2

OpenStudy (aum):

To get a better estimate we can go 0.5 on either side but you have to calculate (0.5)^3 and (1.5)^3 if you want to do that.

OpenStudy (anonymous):

i see. but yea, id rather use whole numbers since i'm not allowed to use a calculator

OpenStudy (aum):

Actually 0.1 on either side may not be that hard to do.

OpenStudy (anonymous):

should there be another equation of v(1) in addition to the one you posted earlier, aum?

OpenStudy (anonymous):

oh

OpenStudy (aum):

\[ v(1) \approx \frac { s(1.1) - s(0.9) }{ 1.1 - 0.9 } = \frac { 1.1^3 + 1.1 - 0.9^3 - 0.9 }{ 0.2} = \\ \text{ } \\ \frac { 1.331 + 1.1 - 0.729 - 0.9 }{ 0.2} = \frac{0.802}{0.2} = \frac{0.802}{0.2} * \frac{5}{5} = \frac{4.01}{1} = 4.01 \]

OpenStudy (anonymous):

thank you to both; you taught me stuff i didn't know

OpenStudy (aum):

You are welcome.